OpenStudy (saifoo.khan):

Restriction? (x+8)/(x^2 +4x-32) i think it's x=4?

OpenStudy (akshay_budhkar):

the denominator is not zero

OpenStudy (karatechopper):

12/0? undefined to me..

OpenStudy (akshay_budhkar):

so u got two answers (as it is a quadratic )

OpenStudy (saifoo.khan):

This is the question. (x+8)/(x^2 +4x-32)

OpenStudy (akshay_budhkar):

(x^2 +4x-32) is not zero

OpenStudy (karatechopper):

restriction=undefined?

OpenStudy (saifoo.khan):

So restriction?

OpenStudy (saifoo.khan):

Idk how to find it's restriction.

OpenStudy (anonymous):

(x^2 +4x-32) = 0 (x + 8) (x-4) = 0 x cannot be equal to -8 or 4

OpenStudy (saifoo.khan):

Oh, okay, so

OpenStudy (saifoo.khan):

Ah, Righto!

OpenStudy (saifoo.khan):

Restriction = \[x \ne -8 , x \ne 4\]

OpenStudy (akshay_budhkar):

yes

OpenStudy (saifoo.khan):

Thankyou!! ^_^

OpenStudy (akshay_budhkar):

yw :D :P